CARA Telecon Minutes Date: May 7, 2001 Attendance: BU: T. Bania Cornell CMU: J. Peterson Northwestern: G. Novak R.I.T: S.A.O.: T. Stark, A. Lane. W. Walsh South Pole: C. Martin, C. Kaminski Chicago: J. Carlstrom, A. Harper, S. Meyer, N. Odalen R. Pernic, R. Landsberg, J. Bausch, Erik Leitch, R. Loewenstein UCB: UCSB: J. Ruhl Approval of previous minutes: It was requested that CARA members review the previous minutes and send changes/corrections and approval via email to John Carlstrom. Agenda Items News from the Pole: C. Martin reported that AST/RO has been down the last 4-5 days with encoder problems. Viper is doing drift scans but there is an intermittent problem with a limit switch. M. Newcomb still does not understand what is causing this problem. AASTO is working well. C. Kaminski has not been knocking off ice for awhile at least since the pressure was fixed. The new Doctor (Betty) is fitting in well. She has more ice time then most people. There is a nurse that has been at Pole since station close. 20% of this individuals time is spent as a nurse and the balance as manager of waste management. AST/RO - T. Stark reported that there are spares at the Pole for the encoder. He believes that the AST/RO team has a handle on the problem. This problem has come up over the years but problem has been intermittent. Suspects that Chris Martin will be able to fix it shortly. Viper - J. Ruhl reported that M. Newcomb is mostly doing drift scans at this point. Did moon door down of space building, as they want to map chopper throw well. Attempted at end of summer (last day) but the scan died. Repeated, but still need to do it again as there was a pointing problem. This is needed to understand the position of the acbar beam on the sky as a function of chopper angle. "Drift scans look stable enough. Limit switch is making us lose some data. We don't know what is going on, but not dead in the water". DASI - going well - plugging away. Additional fans installed to cool correlator crate - not sure if they have been running hotter than normal, but runs cool enough with additional fans. Backup computer re-boots by itself; could be due to dirty power at MAPO. C. Kaminski reported that power would be cut off at Pole for 10-15 seconds today and then again tomorrow. We should check on power. C. Kaminski is having problem with SODAR - it has to be turned back on every few days -- related to DASI computer reboots??? ***C. Kaminski will get line checker and check out possible problem with power to these computers.*** Tipper is plugging away. FTS - cooled down and in operating condition. Working first week of every month. Hopefully window will be open this week as tau was down to 1.0. TEG problem was not due to icing problem. The propane tank dropped pressure when temperature dropped. C. Kaminski put high-pressure regulator on. This brought press up to normal and has been running since. *** R. Pernic to write up short paragraph on problem/fix so we can add to minutes.** (ed. note: see memo circulated) SIPS - R. Pernic reported that he has received comments from AST/RO and Acbar groups only. R. Pernic needs SIP project descriptions from everyone that has either a CARA funded or newly funded OPP project. R. Pernic had good meeting at Raytheon. Not much happening on SIPs until Doctor was returned from Pole. Raytheon has a hard copy of the CARA SIP. Helium - J. Ruhl reported that J. Peterson would go to PA around 5/21/01. R. Pernic should also go. G. Novak will go should either Peterson or Pernic not be able to make the trip. ***J. Ruhl will send email to Raytheon that Peterson/Pernic will make the trip and to contact them in the future.*** Liquid Helium update: (Ruhl) 1000-1200 liters in Wessington. 3000 gallon is boiling off as expected. Best-case estimate based on flow gives 8 months of helium left if just left to boil off (ow working on pressure guage estimate; looks worse than that). The next critical information will come when the 3000 gal dewar is tapped -- thought is that last year's early loss of helium was due to very lossy transfers out of 3000 gal dewar (due to leak in vacuum jacket of fill line on 3000 gal dewar). Status of nitrogen plant unknown. Funding notification sent by NSF to the University of Chicago. Chicago will begin preparing subcontracts. REU - R. Landsberg has funding to support REU students. Contact him with your requests. Post CARA status of funding. DASI funded; cut back a bit. Acbar funded at some level. SPARO declined but G. Novak working on a new proposal and is re-submitting. IR AIRO declined. Only 4 reviews solicited! SPIFI funded at some level. SPIFI and ACBAR also received some money from NSF-OPP Infrastructure grants for upgrades. Gordon please confirm: ** SPIFI will not be going to Pole in 2002 soonest 11/2003 ? ** We need to get together as a group and plan what is needed from Polar Ops. Then go to NSF with plan and request for funds. Need to move on this soon, say by 6/30/01. We will be going to E. Chang and C. Erb with our request as should come out of NSF operations funds not science.